OpenStudy (anonymous):

use integration by parts to compute the following integral, please show steps!!! integrate: sin(3lnx)dx

OpenStudy (anonymous):

\[t=\ln x\] \[x=e^{t}\] \[result=\int\limits_{}^{}\sin t e^{t}dt =\int\limits_{}^{}0.5(\sin t +\cos t+\sin t -\cos t )e^{t}dt\] \[ =0.5 \int \limits_{}^{}\sin t e^{t} + \cos t e^{t}dt+0.5 \int \limits_{}^{}\sin t e^{t} - \cos t e^{t}dt\] \[=0.5 (\sin t e^{t} - \cos t e^{t}) \] \[=0.5 x(\sin \ln x - \cos \ln x) \]

OpenStudy (anonymous):

what happened to the 3?